I need help with this word problem:
A pineapple and 4 oranges cost $8.40. The pineapple costs $5.90 more than an orange. Find the cost of the pineapple.
I did 8.40-5.90=2.50 \ 4= .63. 5.90 + .63 = 6.53.
Pineapple cost $6.53.
Did I do that right?

Answer by DrBeeee(684)

Let o = cost of an orange
Let p = cost of a pineapple
Then
(1) p = o + 590
and
(2) p + 4*o = 840
Substitute (1) into (2) and get
(3) o + 590 + 4*o = 840 or
(4) 5*o = 250 or
(5) o = 50
Answer: Oranges cost 50cts apiece and a pineapple costs $6.40
Your answer is wrong because you didn't use 590 PLUS the cost of an orange. So you divided 250 by 4, not 5 as it should be.
Please don't use "guess-work" to solve algebraic problems, use algebra! Define the variables, write appropriate equations, then solve. It will be easy, believe me. It's not a quiz show, it's math.